# What Is Missing in Multilingual Visual Reasoning and How to Fix It

### v1

@@ Abstract
NLP models today strive for supporting multiple languages and modalities, improving accessibility for diverse users. In this paper, we evaluate their multilingual, multimodal capabilities by testing on a visual reasoning task. We observe that proprietary systems like GPT-4V obtain the best performance on this task now, but open models lag in comparison. Surprisingly, GPT-4V exhibits similar performance between English and other languages, indicating the potential for equitable system development across languages. Our analysis on model failures reveals three key aspects that make this task challenging: multilinguality, complex reasoning, and multimodality. To address these challenges, we propose three targeted interventions including a translate-test approach to tackle multilinguality, a visual programming approach to break down complex reasoning, and a novel method that leverages image captioning to address multimodality. Our interventions achieve the best open performance on this task in a zero-shot setting, boosting open model LLaVA by 13.4%, while also minorly improving GPT-4V’s performance. [^1] The code implementations and prompts can be found at https://github.com/yueqis/Multilingual_Visual_Reasoning.git.
@@
Yueqi Song, Simran Khanuja, Graham Neubig Carnegie Mellon University
@@ 1 Introduction
Figure 1: Our Contributions: First, we evaluate the multilingual visual reasoning capabilities of multimodal models; then, we analyze the challenges where these models are falling short; last but not the least, we propose several interventions to address these challenges.
@@
In this work, we first compare and contrast the multilingual, multicultural capabilities of the proprietary system GPT-4V(ision) Achiam et al. (2023) with a plethora of open models like LLaVA Liu et al. (2023c, a), mBLIP Geigle et al. (2023), CCLM Zeng et al. (2023), using two datasets on the same task of reasoning over texts and pairs of images, NLVR2 (Suhr et al., 2019) and MaRVL Liu et al. (2021). We discuss this setup in more details in §2 and §3. We find that GPT-4V significantly outperforms all open models. One additional unprecedented and surprising result is, as shown in Figure 1, GPT-4V’s consistency in performance across all languages, with performance on some even surpassing that on the NLVR2 dataset in English. In contrast, as we will discuss in §4, most open models still show a significant gap between English and other languages, perhaps due to deficiencies in training data, or due to the well-known “curse of multilinguality”, where smaller models are less adept at processing low-resource languages (Conneau et al., 2020). This begs the question: “how can we take open models, and bring them closer to achieving the exciting language-equitable multimodal reasoning results demonstrated by the opaque (and presumably bigger) GPT-4V?”
@@
Towards this end, we conduct a careful analysis of the results from testing models on the multilingual visual reasoning task and discover that failures can arise from any of the three challenging aspects of the task: multilinguality, reasoning, and multimodality. For multilinguality, we find a significant gap in performance for other languages as compared to English. For reasoning, we find a negative correlation of performance and the compositionality of the statement. For multimodality, we find that models were typically pretrained on single image-text pairs, but haven’t seen pairs of images in pretraining, which may lead to a mismatch between pretraining and evaluation objectives. We will discuss this in more details in Section 5.
@@
Based on our analysis, we design three interventions that address these challenges in section 6. The first simply tackles multilinguality – we translate the MaRVL statements to English. Surprisingly, translation leads to a drop in performance for GPT-4V (which might indicate its advanced multilingual capabilities), but helps improve performance for all open models. Our next intervention tackles both multilinguality and reasoning, by generating programs to reason over the set of images using the translated statements, inspired by Gupta and Kembhavi (2023)’s VisProg method. Our third (and most effective) intervention tackles all three challenges by first captioning images conditioned on the statement, and then reasoning over the captions, rather than the images, using chain-of-thought capabilities of text-modality LLMs Wei et al. (2022). Using this intervention, we obtain state-of-the-art zero-shot performance on the MaRVL dataset for open models, and also slightly improve the performance of GPT-4V itself, as shown in Figure 1.
@@ 2 Dataset Description
LLaVA-13B 60.1 54.8 52.6 50.2 55.3 52.9 53.2 -6.9
@@ 3 Models and Evaluation Protocols
We evaluate various open models, including mBLIP (mt0-xl) Geigle et al. (2023), LLaVA Liu et al. (2023c, a), CCLM Zeng et al. (2023), and UNITERs Chen et al. (2020); and a proprietary model GPT-4V(ision).333gpt-4-vision-preview (https://openai.com/research/gpt-4v-system-card), abbreviated as ”GPT-4V”. We describe these models in Section A. We evaluate them in two settings:
Zero-shot. In this setting, the models are not specifically fine-tuned for the task of multimodal visual reasoning. This setting is academically interesting, as it more generally tests the ability of models to perform reasoning tasks, and the results are more likely to be representative of performance on datasets for which training data is not explicitly available. In addition, it is practically useful in that it can also be applied to LMs that cannot as easily by fine-tuned, such as GPT-4V (due to its closed nature), and LLaVa (due to its relatively large size). We test LLaVA, mBLIP, and GPT-4V in this setting.
Finetuned. In addition, for models that can more easily be finetuned, we finetune them on the English NLVR2 dataset, and test them on both NLVR2 and MaRVL. This represents the realistic setting of adapting a multilingual model to a particular reasoning task using only English data, which is relatively available. We test mBLIP, CCLM-4M, xUNITER, and mUNITER in this setting.
@@ 4 How well do proprietary and open models perform on multilingual visual reasoning?
Which model performs the best? Answer: GPT-4V on MaRVL, and mBLIP (mT0-XL) on English post-fintuning. However, in the zero-shot setting, the proprietary model GPT-4V performs the best across all languages, [^4] We put GPT-4V in the zero-shot category because we evaluate the performance of GPT-4V on NLVR2 and MaRVL without finetuning on the NLVR2 training data. However, we do not know if GPT-4V has seen examples of NLVR2 or MaRVL during pretraining. and open models lag behind. Note that despite GPT-4V’s impressive performance, it still lags behind human performance by 10% to 20% across all languages, emphasizing that this task still is not completely solved.
Which open model performs the best? Answer: mBLIP (mT0-XL), regardless of whether it is finetuned. The other open LLM, LLaVA, is not explicitly trained on multilingual data, so the gap in MaRVL and NLVR2 performance is expected.
@@ 5.1 Multilinguality and Sub-Optimal Cross-Lingual Transfer
In the zero-shot setting, GPT-4V and mBLIP both exhibit equitable performance on MaRVL as with NLVR2. While LLaVa is not expected to perform as well for non-English languages, it has poorer performance than mBLIP on NLVR2. While mBLIP is pretrained on multilingual multimodal data, LLaVA is not specifically trained on multilingual data, but it is generally believed that LLaVA has multilingual abilities as it has seen multilingual data during pretraining Liu et al. (2023c, a).
Overall, models have better visual reasoning capabilities when given English inputs and US/European-centric cultures, while still lagging behind when facing multilingual text and multicultural image inputs.
@@ 5.2 Complex Reasoning
Data points in both NLVR2 and MaRVL require complex reasoning. An example statement from NLVR2 is ”one image includes a silver stylus and a device with a blue keyboard base and an open screen propped up like an easel”, which is semantically diverse, long in length, and has a compositional structure, requiring models to perform compositional and complex reasoning to infer the label.
@@
mBLIP (mT0-XL) 67.3 65.2 66.9 68.7 68.2 67.2 67.2 -0.1
LLaVA-13B 60.1 53.1 53.9 54.1 58.3 54.0 54.7 -5.4
@@ 5.3 Multimodality and Mismatch between Pretraining & Evaluation
Multimodal reasoning is known to be harder than reasoning over text alone Mogadala et al. (2021); Park and Kim (2023). Further, NLVR2 and MaRVL contain two images per instance, along with a statement describing both images, while vision-language models are typically pretrained on a single image-text pair Cao et al. (2020), leading to a mismatch in the input between pretraining and evaluation. This, and the inherent difficulty of aligning image data and text data during the reasoning process make this multimodal task particularly challenging. This motivates us to (1) move from processing a pair of images together to processing each image separately; and (2) break down the two modalities of image and text in the reasoning process, as in §6.3.
@@ 6 How can we address these challenges?
Based on our analysis from the previous section, we now move on to examining whether we can devise methods to further improve multilingual multimodal reasoning abilities, particularly those of open models. We examine three research questions, which we discuss in more details in the following subsections respectively. Figure 4 shows a flow chart visualizing the interventions we perform to address the research questions.
@@ 6.1 Addressing Multilinguality: Translate-Test
In §5.1, we find performance on NLVR2 is much better than that on MaRVL. While both are visual reasoning datasets, MaRVL is multi-cultural and contains statements in 5 diverse languages. Since NLP systems perform significantly better with English data Song et al. (2023), we first simply translate the reasoning statements to English using the Google Translate API Wu et al. (2016). A visualization of the process of translate test can be found in Figure 4.
@@
All prior works, as per our knowledge, have observed a gain in performance post translating to English Liu et al. (2021). Our observation is consistent with prior findings for all models, except GPT-4V(ision). All models except for GPT-4V sees an increase in accuracy after performing translate test; while surprisingly, GPT-4V shows a sharp decrease in performance across all MaRVL languages after translate test. However, this is encouraging, because it speaks for the multilingual capabilities of this model, and indicates that the gains provided by translating to English are lower than the errors made in translating cultural-specific nuances in meaning.
For example, the MaRVL statement ”右图有青绿色的苹果” is translated to ”the picture on the right has turquoise apples”, where ”青绿色” is translated to ”turquoise”. However, the color ”青绿色” means pure green with a little bit cyan in Mandarin Chinese, which is different from ”turquoise”. Given this, GPT-4V reasons correctly when provided the statement in Mandarin, but makes mistakes when given the translated statement.
@@ 6.2 Addressing Multilinguality + Reasoning: Visual Programming
Figure 4 shows the flow of VisProg. For example, given the translated statement: There is no one in the bedroom on the left, and there is someone in the bedroom on the right, the generated visual program is:
@@
For this intervention, we use text-davinci-003555text-davinci-003 is the model that the VisProg authors utilized when running VisProg. as a representative of proprietary LLMs and LLaMA2-70B Touvron et al. (2023) to represent open LLMs. Table 3 shows results to this method. Although this method does not achieve as high accuracy as models evaluated end-to-end in Table 1, this approach provides valuable insights of breaking down complex reasoning into modular modules and utilizing prompts to make use of LLMs’ strong in-context abilities. In addition, this approach, without any additional training, achieves on par performance on MaRVL, as compared to ViLT post-fintuning.
@@
Table 4: Captioning Pipeline Performance across Models. For rows with ”No Intervention” stated in the ”Reasoning” column, we pull over the end-to-end results of that model from Table 1, for the sake of comparison.
@@ 6.3 Addressing Multilinguality + Reasoning + Multimodality: Reasoning with Captions
When analyzing errors for NLVR2, Gupta and Kembhavi (2023) note that 69% of them are caused by the vision module. This might be potentially worse for MaRVL, because open visual modules used in VisProg Kim et al. (2021) are trained on Western-centric datasets like Imagenet Russakovsky et al. (2015). Text-based LLMs, on the other hand, are trained on trillions of tokens, and are known to exhibit cultural awareness, albeit it may be limited Yao et al. (2023). Hence, here we target the last remaining challenge, that of multimodal interaction needed for the reasoning process, by working with image captions instead of images. Concretely, we first caption both images, and use an LLM to reason about the given statement with the two captions, rather than with the two images. Figure 4 shows a flow chart of how this pipeline works.
To make sure the captions capture necessary information needed for reasoning about the statement, as a first step of this intervention we use LLMs to generate targeted instructions based on the statement. For instance, consider the statement ”The picture on the left has several pencils of different colors, and the picture on the right has only one pencil” from MaRVL-zh, the targeted instructions for generating captions are:
Left image - ”Write a short caption describing the number and colors of pencils;”
Right image - ”Write a short caption describing the number of pencils”.
@@
As a second step of this method, we generate captions following the targeted instructions in step 1, using multiple captioning models, including InstructBLIP Liu et al. (2023b), PromptCap Hu et al. (2022), GPT-4V, and LLaVA-13B Liu et al. (2023a). The instructions can point the captioning models to focus on targeted contents in the image. For instance, using the same statement in the first step and the images in Figure 6 as an example, the captions generated by GPT-4V are:
Left image - A pencil case containing a single natural wood colored pencil along with other stationery items.;
Right image - A young mind sketches ideas with a classic green pencil, one among a vast array of tools to capture creativity on paper..
For the last step, we prompt LLMs to reason whether the statement match the pair of captions. For instance, using the same example above, GPT-4 reasons as follows:
Reasoning - The statement is False. This is because the left image is described to include only a single two-toned pencil, which implies there are not several pencils of different colors in the left image. However, the statement claims that there should be several pencils of different colors in the left image. Meanwhile, the right image correctly only shows a single green pencil, but the statement is negated by the inaccuracy of the first part regarding the left image.
Using this approach, visio-linguistic models only need to process one image at a time, rather then processing a pair of images simultaneously. In addition, reasoning is performed only over text, instead of over both image and text modalities.
Table 4 show the performance of this intervention. Performance of InstructBLIP and PromptCap under zero-shot setting is on par with UNITER models post-finetuning. For GPT-4V, the overall performance improved minorly; for LLaVA-13B, performance improved by 10% to 16%, depending on language. On average, our pipeline boost LLaVA’s performance on MaRVL by 13.4%, and improving LLaVA’s performance on NLVR2 by 8%. This captioning intervention improves LLaVA’s performance, achieving the best performance under zero-shot setting (without training on pairs of images).

### v2

@@ Abstract
NLP models today strive for supporting multiple languages and modalities, improving accessibility for diverse users. In this paper, we evaluate their multilingual, multimodal capabilities by testing on a visual reasoning task. We observe that proprietary systems like GPT-4V obtain the best performance on this task now, but open models lag in comparison. Surprisingly, GPT-4V exhibits similar performance between English and other languages, indicating the potential for equitable system development across languages. Our analysis on model failures reveals three key aspects that make this task challenging: multilinguality, complex reasoning, and multimodality. To address these challenges, we propose three targeted interventions including a translate-test approach to tackle multilinguality, a visual programming approach to break down complex reasoning, and a method that leverages image captioning to address multimodality. Our interventions achieve the best open performance on this task in a zero-shot setting, boosting open models LLaVA-v1.5-13B by 13.4%, LLaVA-v1.6-34B by 20.3%, and Qwen-VL by 16.7%, while also minorly improving GPT-4V’s performance. [^1] The code implementations and prompts can be found at https://github.com/yueqis/Multilingual_Visual_Reasoning.
@@
Yueqi Song, Simran Khanuja, Graham Neubig {yueqis,gneubig}@cs.cmu.edu Carnegie Mellon University
@@ 1 Introduction
Figure 1: Our Contributions: First, we evaluate the multilingual visual reasoning abilities of various models; then, we analyze key challenges where models are falling short; lastly, we propose three interventions to address these challenges.
@@
In this work, we first compare and contrast the multilingual, multicultural capabilities of the proprietary systems GPT-4V(ision) Achiam et al. (2023) and Gemini 1.5 Pro Team et al. (2023) with a plethora of open models like LLaVA Liu et al. (2023c, a, 2024), Qwen-VL Bai et al. (2023b), Qwen2-VL Wang et al. (2024), Cambrian Tong et al. (2024), Molmo Deitke et al. (2024), Llama Llama Team, Meta (2024), mBLIP Geigle et al. (2023), CCLM Zeng et al. (2023), using two datasets on the same task of reasoning over texts and pairs of images, NLVR2 (Suhr et al., 2019) and MaRVL Liu et al. (2021). We discuss this setup in more details in §2 and §3. We find that GPT-4V significantly outperforms all open models. One additional unprecedented and surprising result is, as shown in Figure 1, GPT-4V’s consistency in performance across all languages, with performance on some even surpassing that on the NLVR2 dataset in English. In contrast, as we will discuss in §4, most open models still show a significant gap between English and other languages, perhaps due to deficiencies in training data, or due to the well-known “curse of multilinguality”, where smaller models are less adept at processing low-resource languages (Conneau et al., 2020). This begs the question: “how can we take open models, and bring them closer to achieving the exciting language-equitable multimodal reasoning results demonstrated by the opaque (and presumably bigger) GPT-4V?”
@@
Towards this end, we conduct a careful analysis of the results from testing models on the multilingual visual reasoning task and discover that failures can arise from any of the three challenging aspects of the task: multilinguality, reasoning, and multimodality. For multilinguality, we find a significant gap in performance for other languages as compared to English. For reasoning, we find a negative correlation of performance and the compositionality of the statement. For multimodality, we find that models were typically pretrained on single image-text pairs, but haven’t seen pairs of images in pretraining, which may lead to a mismatch between pretraining and evaluation objectives. We will discuss this in more details in §5.
@@
Based on our analysis, we design three interventions that address these challenges in section 6. The first simply tackles multilinguality – we translate the MaRVL statements to English. Surprisingly, translation leads to a drop in performance for GPT-4V and Gemini-1.5-Pro (which might indicate their advanced multilingual capabilities), but helps improve performance for open models. Our next intervention tackles both multilinguality and reasoning, by generating programs to reason over the set of images using the translated statements, inspired by Gupta and Kembhavi (2023)’s VisProg method. Our third (and most effective) intervention tackles all three challenges by first captioning images conditioned on the statement, and then reasoning over the captions, rather than the images, using chain-of-thought capabilities of text-modality LLMs Wei et al. (2022). Using this intervention, we obtain state-of-the-art zero-shot performance on the MaRVL dataset for open models, and also slightly improve the performance of GPT-4V itself, as shown in Figure 1.
@@ 2 Dataset Description
LLaVA-v1.5-13B 60.1 54.8 52.6 50.2 55.3 52.9 53.2 -6.9
LLaVA-v1.6-34B 54.9 56.0 51.8 43.4 57.9 55.3 52.9 -2.0
Qwen-VL 60.3 54.5 50.7 50.3 55.4 58.4 53.9 -6.4
Qwen2-VL-7B-Instruct 81.5 73.5 54.8 60.5 69.9 75.1 66.2 -15.3
Cambrian-8B 75.4 64.7 53.6 56.7 65.2 68.9 61.8 -13.6
Molmo-7B 65.3 61.1 49.6 49.6 52.2 62.2 54.9 -10.4
Llama3.2-11B 64.5 62.7 52.4 54.0 61.6 59.5 58.0 -6.5
@@ 3 Models and Evaluation Protocols
We evaluate various open models, including mBLIP (mt0-xl) Geigle et al. (2023), LLaVA Liu et al. (2023a, 2024), Qwen-VL Bai et al. (2023b), Qwen2-VL-7B-Instruct Wang et al. (2024), Cambrian-8B Tong et al. (2024), Molmo-7B Deitke et al. (2024), CCLM Zeng et al. (2023), and UNITERs Chen et al. (2020); and a proprietary model GPT-4V(ision).333gpt-4-vision-preview (https://openai.com/research/gpt-4v-system-card), abbreviated as ”GPT-4V”. We describe these models in §A. We evaluate them in two settings:
Zero-shot. In this setting, models are not specifically fine-tuned for the task of visual reasoning. This setting is academically interesting, as it more generally tests the ability of models to perform tasks, and the results are more likely to be representative of performance on datasets for which training data is not available. In addition, it is practically useful since it can also be applied to LMs that cannot as easily be fine-tuned, such as the proprietary models GPT-4V and Gemini 1.5 Pro (due to their closed nature), and some large open models such as LLaVA and Qwen-VL (due to their relatively large sizes). We test LLaVA, Qwen-VL, Qwen2-VL-7B-Instruct, Cambrian-8B, Molmo-7B, mBLIP, GPT-4V, and Gemini-1.5-Pro in this setting.
Finetuned. We finetune models that can more easily be finetuned on the English NLVR2 dataset, and test on NLVR2 and MaRVL. This represents the realistic setting, adapting multilingual models to particular tasks using English data, which is relatively available. We test mBLIP, CCLM-4M, xUNITER, and mUNITER in this setting.
@@ 4 How well do proprietary and open models perform on multilingual visual reasoning?
Which model performs the best? Answer: GPT-4V on MaRVL, and mBLIP (mT0-XL) on English post-fintuning. However, in the zero-shot setting, the proprietary model GPT-4V performs the best across all languages other than English, [^4] We put GPT-4V in the zero-shot category because we evaluate the performance of GPT-4V on NLVR2 and MaRVL without finetuning on the NLVR2 training data. However, we do not know if GPT-4V has seen examples of NLVR2 or MaRVL during pretraining. and open models lag behind especially in the multilingual setting. Note that despite GPT-4V’s impressive performance, it still lags behind human performance by 10% to 20% across all languages, emphasizing that this task still is not completely solved.
Which open model performs the best? Answer: mBLIP (mT0-XL), regardless of whether it is finetuned. The other open LMMs, for example LLaVA and Qwen-VL, are not explicitly trained on multilingual data, so the gap in MaRVL and NLVR2 performance is expected.
@@ 5.1 Multilinguality and Sub-Optimal Cross-Lingual Transfer
In the zero-shot setting, GPT-4V and mBLIP both exhibit equitable performance on MaRVL as with NLVR2. Gemini 1.5 Pro also demonstrates equitable performance among languages to some extent. While LLaVA, Cambrian, Molmo, and Llama are not expected to perform as well for non-English languages and Qwen is not expected to perform as well for non-English and non-Chinese languages, they have poorer performance than mBLIP on NLVR2. While mBLIP is pretrained on multilingual multimodal data, LLaVA is not specifically trained on multilingual data. However, Qwen-VL is pretrained on Chinese data Bai et al. (2023b), and it is generally believed that LLaVA has multilingual abilities as it has seen multilingual data during pretraining Liu et al. (2023c, a, 2024).
Overall, models have better visual reasoning abilities when given English inputs from US/European-centric cultures, while still lagging behind when facing multilingual and multicultural inputs.
@@ 5.2 Complex Reasoning
Data points in both NLVR2 and MaRVL require complex reasoning. An example statement from NLVR2 is "one image includes a silver stylus and a device with a blue keyboard base and an open screen propped up like an easel", which is semantically diverse, long in length, and has a compositional structure, requiring models to perform compositional and complex reasoning to infer the label.
@@
Gemini 1.5 Pro 76.4 70.1 65.3 71.0 71.9 72.6 70.2 -6.2
LLaVA-v1.5-13B 60.1 53.1 53.9 54.1 58.3 54.0 54.7 -5.4
LLaVA-v1.6-34B 54.9 55.7 53.1 52.8 55.3 55.4 54.5 -0.4
Qwen-VL 60.3 58.2 56.0 58.8 63.0 58.4 58.9 -1.42
@@ 5.3 Multimodality and Mismatch between Pretraining & Evaluation
NLVR2 and MaRVL contain two images per instance, along with a statement describing them, while vision-language models are typically trained on a single image-text pair Cao et al. (2020), leading to a mismatch in the input between pretraining and evaluation. Further, multimodal reasoning is known to be harder than reasoning over text alone Mogadala et al. (2021); Park and Kim (2023). Although Qwen has seen multi-image inputs during training Bai et al. (2023b), it still encounters difficulties in handling the complexities presented by multimodal reasoning during evaluation.
These, and the inherent difficulty of aligning image data and text data during the reasoning process make this task particularly challenging. This motivates us to (1) move from processing a pair of images together to processing each image separately; and (2) break down the two modalities of image and text in the reasoning process, as in §6.3.
@@ 6 How can we address these challenges?
Based on our analysis from the previous section, we now move on to examining whether we can devise methods to further improve multilingual multimodal reasoning abilities, particularly those of open models. We examine three research questions, which we discuss in more details in the following subsections respectively. We will focus on a subset of the models from Section 3. Figure 4 shows a flow chart visualizing the interventions we perform to address the research questions555§C discusses additional computation cost incurred by the interventions..
@@ 6.1 Addressing Multilinguality: Translate-Test
In §5.1, we find performance on NLVR2 is much better than MaRVL. While both are visual reasoning datasets, MaRVL is multi-cultural and contains data in 5 diverse languages. Since NLP systems perform significantly better with English data Song et al. (2023), we first simply translate the reasoning statements to English using the Google Translate API Wu et al. (2016). A visualization of the process of translate test can be found in Figure 4.
@@
All prior works, as per our knowledge, have observed a gain in performance post translating to English Liu et al. (2021). Our observation is consistent with prior findings for all models, except GPT-4V(ision) and Gemini 1.5 Pro. All models except for GPT-4V and Gemini 1.5 Pro see an increase in accuracy after performing translate test; while surprisingly, GPT-4V and Gemini 1.5 Pro show a sharp decrease in performance across almost all MaRVL languages after translate test. However, this is encouraging, because it speaks for the multilingual capabilities of these models, and indicates that the gains provided by translating to English are lower than the errors made in translating cultural-specific nuances in meaning.
{CJK*}
UTF8gbsn For example, the MaRVL statement "右图有青绿色的苹果" is translated to "the picture on the right has turquoise apples", where "青绿色" is translated to "turquoise". However, the color "青绿色" means pure green with a little bit cyan in Mandarin Chinese, which is different from "turquoise". Given this, GPT-4V reasons correctly when provided the statement in Mandarin, but makes mistakes when given the translated statement [^6] See discussion on whether translate test may introduce biases due to inaccuracies in translation in Appendix D..
@@ 6.2 Addressing Multilinguality + Reasoning: Visual Programming
Figure 4 shows the flow of VisProg. For example, given the statement: There is no one in the bedroom on the left, and there is someone in the bedroom on the right, the generated visual program is:
@@
For this intervention, we use text-davinci-003777text-davinci-003 is the model that the VisProg authors utilized when running VisProg. as a representative of proprietary LLMs and LLaMA2-70B Touvron et al. (2023) to represent open LLMs. Table 3 shows results to this method. Although this method does not achieve as high accuracy as models evaluated end-to-end in Table 1, this approach provides valuable insights of breaking down complex reasoning into modular modules and utilizing prompts to make use of LLMs’ strong in-context abilities. In addition, this approach, without any additional training, achieves on par performance on MaRVL, as compared to ViLT post-fintuning.
@@
LLaVA-v1.6-34B No Intervention 54.9 56.0 51.8 43.4 57.9 55.3 52.9
LLaMA2-70B 77.8 75.9 71.3 71.2 80.6 78.3 75.5
Qwen-VL No Intervention 60.3 54.5 50.7 50.3 55.4 58.4 53.9
LLaMA2-70B 70.3 72.1 66.3 65.1 76.7 72.8 70.6
Table 4: Captioning Pipeline Performance across Models. For rows with "No Intervention" stated in the "Reasoning" column, we pull over the end-to-end results of that model from Table 1, for the sake of comparison.
@@ 6.3 Addressing Multilinguality + Reasoning + Multimodality: Reasoning with Captions
When analyzing errors for NLVR2, Gupta and Kembhavi (2023) note that 69% of them are caused by the vision module. This might be potentially worse for MaRVL, because open visual modules used in VisProg Kim et al. (2021) are trained on Western-centric datasets like Imagenet Russakovsky et al. (2015). Text-based LLMs, on the other hand, are trained on trillions of tokens, and are known to exhibit cultural awareness, albeit it may be limited Yao et al. (2023). Hence, here we target the last remaining challenge, that of multimodal interaction needed for the reasoning process, by working with image captions instead of images. Concretely, we first caption both images, and use LLMs to reason about the statement with the two captions, instead of with the two images. Figure 4 shows a flow chart of how this pipeline works.
To make sure the captions capture necessary information needed for reasoning about the statement, as a first step of this intervention we use LLMs to generate targeted instructions based on the statement. Consider the statement "The picture on the left has several pencils of different colors, and the picture on the right has only one pencil" from MaRVL-zh, the targeted instructions are:
Left image - "Write a short caption describing the number and colors of pencils;"
Right image - "Write a short caption describing the number of pencils".
@@
As a second step, we generate captions following the targeted instructions in step 1, using various captioning models, including InstructBLIP Liu et al. (2023b), PromptCap Hu et al. (2022), GPT-4V, LLaVA-v1.5-13B Liu et al. (2023a), LLaVA-v1.6-34B Liu et al. (2024), and Qwen-VL Bai et al. (2023b). The instructions can point them to focus on targeted contents in the image. For instance, for the statement in step 1 and the images in Figure 6, the captions generated by GPT-4V are:
Left image - A pencil case containing a single natural wood colored pencil.;
Right image - A young mind sketches ideas with a classic green pencil..
Lastly, we prompt LLMs to reason whether the statement match caption pairs. For instance, for the example above, GPT-4 reasons as follows:
Reasoning - The statement is False. This is because the left image is described to include only a single pencil, which implies there are not several pencils of different colors in the left image. However, the statement claims that there should be several pencils of different colors in the left image. Meanwhile, the right image correctly shows a single green pencil, but the statement is negated by the inaccuracy of the part regarding the left image.
Using this approach, visio-linguistic models only need to process one image at a time, instead of processing a pair of images simultaneously. In addition, reasoning is performed only over text, instead of over both image and text modalities.
Table 4 shows the performance of this intervention. Performance of InstructBLIP and PromptCap under zero-shot setting is on par with UNITER models post-finetuning. This intervention improves performance of LLaVA-v1.5-13B by 10% to 16%, LLaVA-v1.6-34B by 18% to 25%, and Qwen-VL888§B discusses additional experiments on Qwen-VL. by 10% to 21% depending on language, while also minorly improving GPT-4V’s performance. On average, our pipeline boosts LLaVA-v1.5-13B’s performance on MaRVL by 13.4%, LLaVA-v1.6-34B’s performance by 20.3%, and Qwen-VL’s performance by 16.7%. This intervention improves performance of LLaVA and Qwen-VL, achieving the best performance under zero-shot setting (without training on reasoning of pairs of images).